How many moles of H20 are there in 50.0 mL, if the density is 0.988 g/mL?!
Answer #1

Moles= 3.2933 moles

As, density= 0.988 g/mL

And Volume= 50 mL

Hence, mass= 50 × 0.988 = 49.4 grams

Now, moles = mass / molar mass

Molar mass of Water = 18 g/mol

Therefore, moles= 49.4/18 moles = 3.2933 moles